# USTC-ICTS/PCFT-24-24

### v1

@@ VI Conclusions
In this paper, we have discussed the complex structures of boson stars, and established the existence of charge-swapping configurations in the presence of gravity. We have shown how the metric field changes the stability and other properties of the charge-swapping configurations, and recovered the Minkowski flat results in suitable limits. Particularly, the quadratic monomial potential, the sextic polynomial potential and the running mass/logarithmic potential are explored with fully nonlinear numerical relativity simulations. For the quadratic and logarithmic case, the complex boson stars are found to be very stable, and we have yet to see them decay in our long-term simulations. The existence of charge-swapping configurations for the quadratic potential is a novelty when coupling the U(1) scalar to gravity, thanks to the gravitational attractions.
In the free scalar case, taking the dipolar boson star for example, we have found that the real and imaginary parts of the scalar field oscillate with slightly different frequencies. This difference determines the main frequency at which the Noether charge of the two components of the system is exchanged. Also, we have found that configurations can avoid gravitational collapse when the initial stars are prepared with sufficiently low masses and separations, as any excess energy is radiated in the initial relaxation stage. In the sextic potential case, we have found evidence of transient existence of highly compact self-gravitating solutions. Thus, a low scalar field density is not a requirement for the existence of these configurations.
@@ Acknowledgements.
We would like to thank Qi-Xin Xie for helpful discussions. SYZ acknowledges support from the National Key R&D Program of China under grant No. 2022YFC220010 and from the National Natural Science Foundation of China under grant No. 12075233 and 12247103.
Appendix A Convergence tests
The charge-swapping configurations found in this work are very long-lived, so convergence tests of the numerical code is very important. First, we have checked the consistency of the full implementation by constructing and evolving isolated stable boson stars that are previously known, calculating the frequency of the scalar field during the simulation and checking whether it coincides with the frequency of the stationary solution for a long period of time, $t\sim 10^{4}$. The coordinates chosen are such that the different metric functions are expected to stay static during the evolution, which we again verify by plotting the global minima and maxima of some of them. To evaluate the consistency of the charge-swapping dynamics, we have compared with the flat spacetime results in Xie et al. (2021) for the polynomial potential case and Hou et al. (2022) for the logarithmic potential. Specifically, we compared the curves of $E$ and found good agreement despite the fact that the implementation is completely different.
@@ Appendix A Convergence tests
In Fig. 22, we show the Noether charge for different sizes of the physical box and spatial resolutions. We see that the medium and high resolutions provide precise results. The same applies to the two bigger box cases, with $x_{\max}^{i}=160$ and 320, but the three box sizes are in good agreement for the amplitudes, and the frequencies of the charge oscillations differ from each other only by a phase. In our simulations of the dipolar free field case, we used the medium resolution and the smaller box size. Finally, in Fig. 23 we show the $L_{2}$-norm of the Hamiltonian constraint Eq. 24. The violation of this constraint begins from small values and increases in the stage of the merger of the two boson stars. The errors reduce by one order of magnitude for the medium and high resolutions.

### v2

@@ VI Conclusions
In this paper, we have discussed the complex structures of boson stars and established the existence of charge-swapping configurations in the presence of gravity. We have shown how the metric field changes the stability and other properties of the charge-swapping configurations and recovered the Minkowski flat results in suitable limits. Particularly, the quadratic monomial potential, the sextic polynomial potential and the running mass/logarithmic potential are explored with fully nonlinear numerical relativity simulations. For the quadratic and logarithmic case, the complex boson stars are found to be very stable, and we have yet to see them decay in our long-term simulations. The existence of charge-swapping configurations for the quadratic potential is a novelty when coupling the U(1) scalar to gravity, thanks to the gravitational attractions.
In the free scalar case, taking the dipolar boson star for example, we have found that the real and imaginary parts of the scalar field oscillate with slightly different frequencies. This difference determines the main frequency at which the Noether charge of the two components of the system is exchanged. Also, we have found that configurations can avoid gravitational collapse when the initial stars are prepared with sufficiently low masses and separations, as any excess energy is radiated in the initial relaxation stage. In the sextic potential case, we have found evidence of transient existence of compact self-gravitating charge-swapping configurations. Although defining compactness in the case of dynamical configurations is not straightforward, we found that charge-swapping configurations prepared with boson stars with $C\sim 0.1$, some transient configurations emerge after losing about half of their mass (Fig. 13). These resulting structures maintain roughly the same radius as the initial stars, rendering the newly formed configurations compact. Thus, a low scalar field density is not a requirement for the existence of these configurations.
@@ Acknowledgements.
We would like to thank Qi-Xin Xie for helpful discussions. SYZ acknowledges support from the National Key R&D Program of China under grant No. 2022YFC2204603 and from the National Natural Science Foundation of China under grant No. 12075233 and 12247103.
Appendix A Appendix: Convergence tests
The charge-swapping configurations found in this work are very long-lived, so convergence tests of the numerical code are very important. First, we have checked the consistency of the full implementation by constructing and evolving isolated stable boson stars that are previously known, calculating the frequency of the scalar field during the simulation and checking whether it coincides with the frequency of the stationary solution for a long period of time, $t\sim 10^{4}$. The coordinates chosen are such that the different metric functions are expected to stay static during the evolution, which we again verify by plotting the global minima and maxima of some of them. To evaluate the consistency of the charge-swapping dynamics, we have compared with the flat spacetime results in Xie et al. (2021) for the polynomial potential case and Hou et al. (2022) for the logarithmic potential. Specifically, we compared the curves of $E$ and found good agreement despite the fact that the implementation is completely different.
@@ Appendix A Appendix: Convergence tests
In Fig. 22, we show the Noether charge for different sizes of the physical box and spatial resolutions. In this plot we have plotted the time window between $t=3000$ and $t=3400$ to make convergence readily apparent, however the following conclusions can be obtained from other time windows. We see that the medium and high resolutions provide precise results. The same applies to the two bigger box cases, with $x_{\max}^{i}=160$ and 320, but the three box sizes are in good agreement for the amplitudes, and the frequencies of the charge oscillations differ from each other only by a phase. In our simulations of the dipolar free field case, we used the medium resolution and the smaller box size. In Fig. 23 we show the $L_{2}$-norm of the Hamiltonian constraint Eq. 24. The violation of this constraint begins from small values and increases in the initial relaxation stage. The errors reduce by one order of magnitude for the medium and high resolutions.
@@
If we start with different values of $d$ or start with different models of initial boson stars the violation of the constrains (at least in the medium resolution) is of the same order as the one showed in Fig. 23 whenever the configuration do not collapse. For instance, for smaller separations, the constraint violations remain of the same order within a certain range of closely spaced configurations, provided they do not collapse. However, in configurations with initially larger separations, the quantities $H$ and $\mathcal{M}_{i}$ become more pronounced during the initial relaxation stage.
Finally we comment on the momentum constraint. We have tracked the evolution of it and found that the components $L_{2}(\mathcal{M}_{x})$ and $L_{2}(\mathcal{M}_{y})$ exhibit the same qualitative behavior as the Hamiltonian evolution shown in Fig. 23. The maximum occurs at the beggining of the initial relaxation process, followed by some dissipation of the numerical error. Both components are an order of magnitude smaller than the Hamiltonian and the $L_{2}(\mathcal{M}_{z})$ component is approximately half the size of the $x$ and $y$ components.